Using the tap it takes aaron 4 minutes to fill his wading pool. Using the neighbors hose it takes 12 minutes. how long would it take if he used both the tap and the hose?
Found 2 solutions by jorel1380, mananth:
Answer by jorel1380(3719)

t/4+t/12=1
3t+4t=12
7t=12
t=12/7 mins.

Answer by mananth(16949)

Filling the pool 1 job
Tap 4minutes
It does 1/ 4 jobs in 1 minute
Hose 12 minutes
It does 1/12 of thejob in 1 minute
Pipe III 0 minutes
It does #DIV/0! of the job in 1 minute

Together they will do 1/ 4+ 1/12
Together they will do 1/ 3of the job in one minute 0.33
So they will take 3 minutes
m.ananth@hotmail.ca